OpenStudy (anonymous):

there are hundred students in hostel .food for all is for 20 days.if 25 more students join them ,then food last for how many days

OpenStudy (anonymous):

if there are 100 students and food for 20 days - you have to find the ratio - so it'll be: number of students/number of days. this gives: 100/20 = 5. so the food lasts for 5 days. now you have 25 more students. so the total number of students is 125.

OpenStudy (anonymous):

Isn't this an inverse proportion ?

OpenStudy (anonymous):

100/125 = x/20 (cross multiply) 125x = (20)(100) 125x = 2000 x = 2000/125 x = 16 I think the answer is 16 days. I think this is an inverse proportion. I really get confused with these.